find the smallest positive in which lies the following tanx+tanhx=0
1
Expert's answer
2020-09-08T15:02:46-0400

We have to find least positive integer interval in which the solution of tanx+tanhx=0 lies.

Let "f(x) = tan(x) + tanh(x)" .

Now, "f(1) = tan(1) + tanh(1) = 2.319 > 0" ,

"f(2) = tan(2) + tanh(2) = -1.221 < 0" ,

and "f(3) = tan(3)+tanh(3) = 0.8525 > 0" .

We known that "tanh(x) \in (-1,1)" for all values of "x" .

Also, "tan(x)" has infinite discontinuity at "x = \frac{\pi}{2} = 1.57, \frac{3\pi}{2} = 4.71, ..." .

Hence, least positive integer interval in which the solution of tanx+tanhx=0 lies is "(2,3)" .